Show that the function f:R2 to R defined by f(x,y)= x2+y2 is not uniformly continuous


1
Expert's answer
2021-03-24T14:04:13-0400

Inorder to show that f is uniformly continuous, we show that given "\epsilon >0 \exist \, \delta>0 \ni |(x_1,y_1)-(x_2,y_2)| < \delta \implies |f(x_1,y_1) -f(x_2,y_2)|<\epsilon"

Suppose that this is true and consider two points in "\mathbb{R^2}" say (a,b) and "(a + \frac{\delta}{2}, b + \frac{\delta}{2})"

Now consider

"|f(a + \frac{\delta}{2}, b + \frac{\delta}{2})-f(a,b)| = | (a + \frac{\delta}{2})^2 +(b+ \frac{\delta}{2})^2 - (a^2+b^2)| = |(a+b) \delta + \frac{\delta^2}{2}|"

And as "a+b" gets larger, the quantity above is greater than "\epsilon" which is sufficiently small.

So, we have arrived at a contradiction. Hence, we can conclude that f is not uniformly continuous.